OpenStudy (anonymous):

How could you eliminate the fractions to make solving the equation five twenty-thirds m plus twelve twenty-thirds equals twenty-one twenty-thirds easier? Divide both sides of the equation by five twenty-thirds. Divide both sides of the equation by twenty-three. Multiply both sides of the equation by twenty-three. Subtract twelve twenty-thirds from both sides of the equation.

OpenStudy (anonymous):

@DoYourHomework help?

OpenStudy (anonymous):

\[5\frac{ 20 }{ 3 }m + 12\frac{ 20 }{ 3 } = 21\frac{ 20 }{ 3 }?\]

OpenStudy (anonymous):

subtract twelve twenty-thirds from both sides. This will give you \[5\frac{ 20 }{ 3 }m = 9\]

OpenStudy (anonymous):

@DoYourHomework so D. ty